OpenStudy (anonymous):

Put 0.367, 36.5 percent, and 9/25 in order from least to greatest.

OpenStudy (anonymous):

36.5 % = 36.5/100 = 0.365 9/25= 0.36= 0.360 On comparing, 0.36<0.365<0.367 9/25<36.5%<0.367

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

your welcome :)